# Exact Random Graph Matching with Multiple Graphs

### v1

@@ C.1 Proof of Lemma 14
$\leq\binom{n-1}{n^{\delta}}\cdot\mathbb{P}\left(\mathsf{Bin}\left(n^{1+\delta},\alpha\log(n)/n\right)\leq rn^{\delta}\right)$
$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\left(\frac{ne}{n^{\delta}}\right)^{n^{\delta}}\times\left(\frac{\exp\left(r/(\alpha\log n)-1\right)}{(r/(\alpha\log n))^{r/(\alpha\log n)}}\right)^{n^{\delta}\alpha\log(n)}$
$=o(1/n),$
whenever $\delta>1-\alpha$ as desired. Note that (a) uses the Binomial concentration inequality (7) and the fact that $\binom{n-1}{k}\leq\binom{n}{k}\leq\left(\frac{ne}{k}\right)^{k}$. ∎
Our objective is to show that the $k$-core of $G-v$ is sufficiently large with probability $1-o(1/n)$. To that end, consider Algorithm 3 to identify a subset of the $k$-core, originally proposed by Łuczak [Łuc91].
require :  Graph $G$, Set $U\subseteq V(G)$.
1 $U_{0}\leftarrow U$
2 for $i=0,1,2,3,\cdots$ do
4             $U_{i+1}\leftarrow U_{i}\cup\{u\}$
5
6       else
7             return $U_{i}$
8
9       end if
10
11 end for
Algorithm 3 Łuczak expansion
Note that the for loop eventually terminates - the set $V\setminus U_{i}$ is empty, for example when $i=n$ for any input set $U$. The key is to realize that the for loop terminates much faster when the input $U=Z_{k+1}$ , i.e the set of vertices of the input graph $G$ whose degree is $k+1$ or less. Furthermore, the complement of the set output by the algorithm is contained in the $k$-core. Formally,
Lemma 17.
Let $U_{f}$ be the output of Algorithm 3 with input graph $G-v$ and set $U=Z_{k+1}$. Then,
  1. (a)
    $U_{f}^{c}\subseteq\mathsf{core}_{k}(G-v)$.
  2. (b)
    For any $\delta>1-\alpha$,
    $\mathbb{P}\left(|U_{f}|>3n^{\delta}\right)=o(1/n).$
Proof.
(a) The proof is by construction: Since $U_{f}$ is obtained by adding exactly $f$ nodes to $U_{0}$, it follows that $U_{f}^{c}\subseteq U_{0}^{c}=Z_{k+1}^{c}$, so each node in $U_{f}^{c}$ has degree $k+2$ or more in $G-v$. Further, each node in $U_{f}^{c}$ has at most $2$ neighbors in $U_{f}$, else the for loop would not have terminated. Thus, the subgraph of $G-v$ induced on the set $U_{f}^{c}$ has minimum degree at least $k$, and the result follows.
(b) If $|U_{f}|>3n^{\delta}$, then either $|U_{0}|>3n^{\delta}$ or there is some $M$ in $\{0,1,\cdots,f\}$ for which $|U_{M}|=3n^{\delta}$. Therefore,
$\mathbb{P}\left(|U_{f}|>3n^{\delta}\right)$ $\leq\mathbb{P}\left(|U_{0}|>3n^{\delta}\right)+\mathbb{P}\left(\exists\ M\in\{0,1,\cdots,f\}\text{ s.t. }|U_{M}|=3n^{\delta}\right)$
$=o(1/n)+\underbrace{\mathbb{P}\left(\exists\ M\in\{0,1,\cdots,f\}\text{ s.t. }|U_{M}|=3n^{\delta}\right)}_{(\star)},$
by Proposition 16. Note that each iteration $i=0,1,\cdots,M-1$ of the for loop adds exactly $1$ vertex and at least $3$ edges to the subgraph of $G-v$ induced on $U_{M}$. Therefore, the induced subgraph $G|_{U_{M}}$ has $3n^{\delta}$ vertices and at least $3\left(|U_{M}|-|U_{0}|\right)$ edges. Thus,
(($\star$)( ⋆ ) ( $\leq\mathbb{P}\left(\exists\text{ subgraph $H\!=\!(W,F)$ of }G-v\text{ s.t. }|W|=3n^{\delta}\text{ and }|F|\geq 3\left(3n^{\delta}-|U_{0}|\right)\right)$
$\leq\mathbb{P}\left(|U_{0}|>n^{\delta}\right)+\mathbb{P}\left(\exists\text{ subgraph $H\!=\!(W,F)$ of }G-v\text{ s.t. }|W|=3n^{\delta}\text{and }|F|\geq 6n^{\delta}\right)$
$\leq o(1/n)+\underbrace{\binom{n}{3n^{\delta}}\cdot\mathbb{P}\left(\mathsf{Bin}\left(\binom{3n^{\delta}}{2},\frac{\alpha\log(n)}{n}\right)>6n^{\delta}\right)}_{(\star\star)},$
where the last step uses Proposition 16 and a union bound over all possible choices of $W$. Finally, using the relation $\binom{n}{k}\leq\left(\frac{ne}{k}\right)^{k}$ and the concentration inequality (5) from Lemma 13 yields
[…]
$\leq(n^{1-\delta})^{3n^{\delta}}\times\left(\frac{3\alpha e\log n}{4n^{1-\delta}}\right)^{6n^{\delta}}$
$=\left(\frac{3\alpha e\log n}{4n^{(1-\delta)/2}}\right)^{6n^{\delta}}$
$=o(1/n),$
whenever $10<\delta<$ 1. The result follows. ∎
Finally, notice that Lemma 17 directly implies Lemma 14.
Appendix D Proof of Theorem 9
See 9
Proof.
For any vertex cut $U$,
$\left\{\widetilde{c}_{v}(U)>\frac{m^{2}}{4}\left(k+\frac{1}{Cs^{2}}\right)\right\}$ $\stackrel{{\scriptstyle\text{(a)}}}{{\subseteq}}\left\{\widetilde{c}_{v}(U)>|U|\left(m-|U|\right)\left(k+\frac{1}{Cs^{2}}\right)\right\}$
=$=\bigg{\{}\sum_{i\in U}\sum_{j\in U^{c}}\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v)>|U|(m-|U|)\bigg{(}k+\frac{1}{Cs^{2}}\bigg{)}\bigg{\}}$
$\subseteq\bigcup_{i\in U}\bigcup_{j\in U^{c}}\left\{\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v)>k+\frac{1}{Cs^{2}}\right\},$
where (a) uses the fact that the maximum of a set of a numbers is greater than or equal to the average. On the other hand
$\left\{c_{v}(U)=0\right\}=\bigcap_{i\in U}\bigcap_{j\in U^{c}}\left\{v\notin\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})\right\}.$
$p_{1}\leq\sum_{i\in U}\sum_{j\in U^{c}}\mathbb{P}\left(\left\{v\notin\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})\right\}\cap\left\{\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v)>k+\frac{1}{Cs^{2}}\right\}\right)=o(1/n),$
since for any choice of $i$ and $j$, the graph $G_{i}^{\prime}\wedge G_{j}^{\prime}\sim\mathsf{ER}\left(n,Cs^{2}\log(n)/n\right)$. ∎
Appendix E On Stochastic Dominance: Proof of Theorem 10
The objective of this section is to build up to a proof of Theorem 10. We start by making a simple observation about products of Binomial random variables.
Lemma 18.
Let $X_{1},\cdots,X_{m}\sim\mathsf{Bern}(s)$ be i.i.d. random variables, and let $B=X_{1}+\cdots+X_{m}$ denote their sum. For each $\ell$ in $\{1,2,\cdots,\lfloor m/2\rfloor\}$, define
@@ Appendix E On Stochastic Dominance: Proof of Theorem 10
whenever $m-\ell\geq\ell-1$, i.e. $\ell\leq\lfloor m/2\rfloor$. Here, (c) uses the identity $\binom{n}{k}=\binom{n-1}{k-1}+\binom{n-1}{k}$, and the fact that $\binom{n_{1}}{k}\geq\binom{n_{2}}{k}$ whenever $n_{1}\geq n_{2}$. This concludes the proof. ∎
Corollary 19.
Let $F$ be a collection of edges in the parent graph $G$. For any edge $e_{r}\in F$, let $X^{r}_{i}$ denote the indicator random variable $G_{i}^{\prime}(e_{r})\sim\mathsf{Bern}(ps)$. For each $\ell$ in $\{1,\cdots,\lfloor m/2\rfloor\}$, define
@@
Proof.
It suffices to show that $T^{r}_{\ell_{1}}\preceq T^{r}_{\ell_{2}}$ for each $r$, since the edges are independent. Indeed, we have for any $t$ that
@@
which concludes the proof. ∎
With this, we are ready to prove Theorem 10. The theorem is restated for convenience.
See 10
Proof.
Let $\ell_{1},\ell_{2}\in\{1,\cdots,\lfloor m/2\rfloor\}$ such that $\ell_{1}<\ell_{2}$. Let $t\in\mathbb{R}$. Consider the parent graph $G$ and label the set of incident edges on $v$ as $\{e_{1},\cdots,e_{\delta_{G}(v)}\}$. Denote by $X_{i}^{r}$ the indicator random variable $G_{i}^{\prime}(e_{r})\sim\mathsf{Bern}(ps)$. It follows that
@@
as desired. Here, (a) uses Corollary 19. ∎
Appendix F On Low Degree Nodes: Proof of Theorem 11
See 11
Proof.
Consider fixed integers $r_{1},\cdots,r_{m}$ such that $0\leq r_{2},\cdots,r_{m}\leq r$. Since $r$ is constant, by a union bound argument it suffices to show
()=:($(\star)=:\mathbb{P}\left(\left\{\delta_{G_{1}\wedge G_{2}^{\prime}}(v)=r_{2}\right\}\cap\cdots\cap\left\{\delta_{G_{1}\wedge G_{m}^{\prime}}(v)=r_{m}\right\}\right)=o\left(1/n\right).$
Proceed by conditioning on the degree of $v$ in $G_{1}$, which follows a $\mathsf{Bin}(n,ps)$ distribution. Since the degrees of $v$ in the intersection graphs $\{G_{1}\wedge G_{i}^{\prime}:i=2,\cdots,m\}$ are conditionally independent given the degree of $v$ in $G_{1}$, we have
(($\star$)( ⋆ ) $=\mathbb{E}_{D}\left[\mathbb{P}\Big{(}\bigcap_{i=2}^{m}\left\{\delta_{G_{1}\wedge G_{i}^{\prime}}(v)=r_{i}\right\}\Big{|}\delta_{G_{1}}(v)=D\Big{)}\right]$
$=\mathbb{E}_{D}\left[\prod_{i=2}^{m}\mathbb{P}\left(\left\{\delta_{G_{1}\wedge G_{i}^{\prime}}(v)=r_{i}\right\}\Big{|}\delta_{G_{1}}(v)=D\right)\right]$
$=\mathbb{E}_{D}\left[\prod_{i=2}^{m}\binom{D}{r_{i}}s^{r_{i}}(1-s)^{D-r_{i}}\right].$ (13)
Using the fact that $\binom{D}{r_{i}}\leq\left(\frac{De}{r_{i}}\right)^{r_{i}}$, it follows that
@@ Appendix F On Low Degree Nodes: Proof of Theorem 11
[…]
Expanding out the expectation yields
[…]
Proceed by splitting the summation at $(\log n)^{2}$. The first part can be bounded as
$\sum_{d=0}^{\left(\log n\right)^{2}}L_{d}$ $\leq\left(\log n\right)^{2\sum_{i=2}^{m}r_{i}}\sum_{d=0}^{(\log n)^{2}}\left(1-s\right)^{(m-1)d}\cdot\mathbb{P}\left(\mathsf{Bin}(n,ps)=d\right)$
$\leq\left(\log n\right)^{2\sum_{i=2}^{m}r_{i}}\cdot\sum_{d=0}^{n}\left(1-s\right)^{(m-1)d}\cdot\mathbb{P}\left(\mathsf{Bin}(n,ps)=d\right)$
$=\left(\log n\right)^{2\sum_{i=2}^{m}r_{i}}\cdot\mathbb{E}_{D}\left[(1-s)^{(m-1)D}\right]$
=$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\left(\log n\right)^{2\sum_{i=2}^{m}r_{i}}\cdot\left(1-\frac{Cs\left(1-(1-s)^{m-1}\right)\log n}{n}\right)^{n}$
$=o(1/n),$
whenever $Cs(1-(1-s))^{m-1}>1$. Here, (a) is obtained by evaluating the probability generating function of the $\mathsf{Bin}(n,ps)$ random variable at $(1-s)^{m-1}$ and setting $p=C\log(n)/n$.
The other part of the sum can now be bounded as follows.
$\sum_{d=\left(\log n\right)^{2}}^{n}L_{d}$ $\leq\left[\max_{d\colon(\log n)^{2}\leq d\leq n}d^{\sum_{i=2}^{m}r_{i}}(1-s)^{md}\right]\cdot\mathbb{P}\left(\mathsf{Bin}(n,ps)\geq(\log n)^{2}\right)$
$\stackrel{{\scriptstyle\text{(b)}}}{{\leq}}\left[\left(\log n\right)^{2\sum_{i=2}^{m}r_{i}}(1-s)^{m\left(\log n\right)^{2}}\right]\times 2^{-\left(\log n\right)^{2}}$
$=(\log n)^{2\sum_{i=2}^{m}r_{i}}\left(\frac{(1-s)^{m}}{2}\right)^{(\log n)^{2}}$
$=o\left(1/n\right).$
whenever $C>0$. Here, (b) is true because the function $d\mapsto d^{\sum r_{i}}(a-s)^{md}$ is decreasing on the interval $[(\log n)^{2},n]$ for all sufficiently large $n$. Finally, the concentration inequality for the Binomial distribution holds by (6) in Lemma 13. The inequality applies since $p=C\log(n)/n$ and since $(\log n)^{2}>6Cs\log(n)$ for all $n$ sufficiently large. This concludes the proof. ∎

### v2

@@ B.1 Proof of Lemma 9
Appendix C Stochastic domination of crossing edges in ~$\widetilde{\mathcal{H}}(v)$: Proof of Lemma 25
The objective of this section is to build up to a proof of Lemma 25. We start by making a simple observation about products of Binomial random variables.
Lemma 30
Let $X_{1},\cdots,X_{m}\sim\mathrm{Bern}(s)$ be i.i.d. random variables, and let $B=X_{1}+\cdots+X_{m}$ denote their sum. For each $\ell$ in $\{1,2,\cdots,\lfloor m/2\rfloor\}$, define
@@ Appendix C Stochastic domination of crossing edges in ~$\widetilde{\mathcal{H}}(v)$: Proof of Lemma 25
whenever $m-\ell\geq\ell-1$, i.e. $\ell\leq\lfloor m/2\rfloor$. Here, (c) uses the identity $\binom{n}{k}=\binom{n-1}{k-1}+\binom{n-1}{k}$, and the fact that $\binom{n_{1}}{k}\geq\binom{n_{2}}{k}$ whenever $n_{1}\geq n_{2}$. This concludes the proof. $\Box$
Corollary 31
Let $F$ be a collection of edges in the parent graph $G$. For any edge $e_{r}\in F$, let $X^{r}_{i}$ denote the indicator random variable $G_{i}^{\prime}(e_{r})\sim\mathrm{Bern}(ps)$. For each $\ell$ in $\{1,\cdots,\lfloor m/2\rfloor\}$, define
@@
Proof. It suffices to show that $T^{r}_{\ell_{1}}\preceq T^{r}_{\ell_{2}}$ for each $r$, since the edges are independent. For any $t$,
@@
[…]
With this, we are ready to prove Lemma 25. The lemma is restated for convenience.
See 25 Proof. Let $\ell_{1},\ell_{2}\in\{1,\cdots,\lfloor m/2\rfloor\}$ such that $\ell_{1}<\ell_{2}$. Let $t\in\mathbb{R}$. Consider the parent graph $G$ and label the set of incident edges on $v$ as $\{e_{1},\cdots,e_{\delta_{G}(v)}\}$. Denote by $X_{i}^{r}$ the indicator random variable $G_{i}^{\prime}(e_{r})\sim\mathrm{Bern}(ps)$. It follows that